# Parameter estimation for Gibbs distributions

### v9

@@ 8.2 Bounds for ${P_{\tt count-weak}}$ in the general integer setting
Define $m=\lfloor n/2\rfloor$. We have $d=m$ instances, where we set $\beta_{\min}=0,\beta_{\max}=m\log 2$. We define $c$ by:
@@
Let us provide some intuition about this instance. For each $i=1,\dots,m$, there is a small count $c_{2i-1}$ hidden between two larger counts $c_{2i-2},c_{2i}$. The only way to determine the precise value of $c_{2i-1}$ is to take roughly $\frac{1}{\delta\varepsilon^{2}}$ samples of $\mu_{\beta}$ where $\beta\approx i\log 2$. If $\beta\gg i\log 2$, then $\mu_{\beta}(2i)$ will drown out $\mu_{\beta}(2i-1)$, and likewise if $\beta\ll i\log 2$ then $\mu_{\beta}(2i-2)$ will drown out $\mu_{\beta}(2i-1)$. Consequently, to solve ${P^{\delta,\varepsilon}_{\tt count}}$, we need roughly $\Omega(\frac{m}{\delta\varepsilon^{2}})$ samples in total.
@@
Suppose $\delta<1/100$. Then for any integer $k\in\mathbb{Z}$, we have $Z(k\log 2)\leq 2^{k^{2}+2}$.
@@
For $n\geq n_{0},q>q_{0},\delta<1/100,\varepsilon<1/100$, where $n_{0},q_{0},\delta_{0}$ are absolute constants, the instance can be constructed so that ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{\min\{\sqrt{q},n\}\log(1/\gamma)}{\delta\varepsilon^{2}})$.
@@
By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, it suffices to show this under the assumption that $n<\sqrt{q}/10$. In this case, we apply Corollary 46(b). Here $\kappa=3\varepsilon$. Thus, $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}/4+6\varepsilon\leq q$ for $q_{0}$ sufficiently large.
For each instance $i$, we use witnesses $x=0,y=2i-1$, where $\Delta(x),\Delta(y)\geq 2\delta\geq\delta e^{2\kappa}$ for $\delta<1/100$, and $|\lambda_{x}^{(i)}-\lambda_{y}^{(i)}|=3\varepsilon$. So ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})\ =\Omega\bigl{(}\frac{n\log(1/\gamma)}{\delta\varepsilon^{2}}\bigr{)}$. ∎
@@ 8.3 Bounds for the continuous setting
In the constructions of Sections 8.1 and 8.2 for the integer setting, the complexity bounds have a term $\min\{q,n^{2}\}$. In the continous setting, this dependence on $n$ can be removed; we can get lower bounds which depend only on $q$. Namely, given some family $c^{(i)}$ of integer-setting distributions (supported on $\{0,1,\dots,n\})$, we can construct a new continous-setting instance via:
$\tilde{c}_{x}=c_{n(x-1)},\tilde{\lambda}^{(i)}_{x}=\lambda^{(i)}_{n(x-1)}\text{ for all $x,i$},\qquad\qquad\tilde{\beta}_{\min}=\beta_{\min}n,\tilde{\beta}_{\max}=\beta_{\max}n.$
Crucially here, the new instance $\tilde{c}$ is supported on $x\in[1,2]$, i.e. it has parameter $\tilde{n}=2$. Furthermore observe that
$\tilde{Z}^{(i)}(\beta)=\sum_{x}\tilde{c}_{x}^{(i)}e^{\beta x}=\sum_{x}c^{(i)}_{n(x-1)}e^{\beta x}=\sum_{y}c_{y}^{(i)}e^{\beta(1+y/n)}=e^{\beta}Z^{(i)}(\beta/n).$
So $\tilde{Q}^{(i)}(\tilde{\beta}_{\max})=e^{n(\beta_{\max}-\beta_{\min})}Q(\beta_{\max})$, and $\tilde{\mu}_{\beta n}^{(i)}(1+x/n)=\mu_{\beta}^{(i)}(x)$. Thus, solving ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ on the instances $\tilde{c}$ is equivalent to solving them on the original instances $c$. Furthermore, in order to use Corollary 46 on the new problem instances $\tilde{c}$, the original instances $c$ must satisfy
$n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq q$
To get the lower bound $\Omega(\frac{q}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ and ${P^{\delta,\varepsilon}_{\tt count-weak}}$ in the continuous setting, we construct the problem instances $c$ from Section 8.1 for value $n=\lfloor\sqrt{q}/10\rfloor$. Lemma 50 gives $n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}+n^{2}+10\varepsilon$, which is at most $q$ for $q\geq q_{0}$ sufficiently large.
Similarly, for the lower bound $\Omega(\frac{\sqrt{q}}{\delta\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ in the continous setting, we construct the problem instances $c$ of Section 8.2 for value $n=\lfloor\sqrt{q}/10\rfloor$. For $m=\lfloor n/2\rfloor$, we have $n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq n(m\log 2-0)+n^{2}/4+3\varepsilon$, which is at most $q$ for $q\geq q_{0}$ sufficiently large.
@@ Appendix A Proofs for error bounds of ${P^{\delta,\varepsilon}_{\tt count}}$
Proof of Theorem 5. Note that $\mu_{\alpha}(x)=\pi(x)e^{\alpha x}/Q(\alpha)$. If $\pi(x)=0$, then the approximation guarantee of ${P^{\delta,\varepsilon}_{\tt count}}$ directly gives $\hat{\mu}_{\alpha}(x)=0=\mu_{\alpha}(x)$. Otherwise, we can use the approximation guarantees of ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$ to show the upper bound on $\hat{\mu}_{\alpha}(x)$:
@@
$\leq e^{\varepsilon}\mu_{\alpha}(x)(1+\varepsilon(1+\delta/\mu_{\alpha}(x)))\qquad\qquad\qquad\qquad\qquad\qquad(\text{$\mu_{\alpha}(x)\leq\Delta(x)$ by definition})$
$=e^{\varepsilon}(1+\varepsilon)\mu_{\alpha}(x)+e^{\varepsilon}\delta\leq\mu_{\alpha}(x)+3\varepsilon(\mu_{\alpha}(x)+\delta)\qquad\qquad\ (\text{straightforward calculus for $\varepsilon<1/2$})$
@@ Appendix B Correctness with approximate oracles
Here, we define the total variation distance $||\cdot||_{TV}$ by
$||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}=\max_{{\cal K}\subseteq\mathcal{F}}|\tilde{\mu}_{\beta}({\cal K})-\mu_{\beta}({\cal K})|=\tfrac{1}{2}\sum_{x}|\tilde{\mu}_{\beta}(x)-\mu_{\beta}(x)|$
@@
Suppose that sampling procedure $\mathfrak{A}$ has cost at most $T$ for some known parameter $T$. Let $\tilde{\mathfrak{A}}$ be the algorithm obtained from $\mathfrak{A}$ as follows: (i) we replace calls $x\sim\mu_{\beta}$ with calls $x\sim\tilde{\mu}_{\beta}$ where $||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}\leq\gamma/T$; (ii) we terminate algorithm after $T/\gamma$ steps and return arbitrary answer.
Then $\tilde{\mathfrak{A}}$ has cost $O(T)$, and the total variation distance of the of $\mathfrak{A}$ and $\mathfrak{A}^{\prime}$ is at most $2\gamma$.
@@
The distributions $\mu_{\beta}$ and $\tilde{\mu}_{\beta}$ can be maximally coupled so that samples $x\sim\mu_{\beta}$ and $x\sim\tilde{\mu}_{\beta}$ are identical with probability at least $1-\gamma/T$. Assume the $k^{\text{th}}$ call to $\mu_{\beta}$ in $\mathfrak{A}$ is coupled with the $k^{\text{th}}$ call to $\tilde{\mu}_{\tilde{\beta}}$ in $\tilde{\mathfrak{A}}$ when both calls are defined and $\beta=\tilde{\beta}$. We say the $k^{\text{th}}$ call is good if either (i) both calls are defined and the produced samples are identical, or (ii) $\mathfrak{A}$ has terminated earlier. Note that ${\mathbb{P}}[\mbox{$k^{\text{th}}$ call is good}\>|\>\mbox{all previous calls were good}]\geq 1-\gamma/T$, since the conditioning event implies $\beta=\tilde{\beta}$ (assuming the calls are defined).
@@ 2. $\mathcal{E}_{2}$:
    The number of oracle calls by $\mathfrak{A}$ is at most $T/\gamma$. By Markov’s inequality, this has probability at least $1-\frac{{\mathbb{E}}[A]}{T/\gamma}\geq 1-\gamma$.
If these two events occur, then $\tilde{\mathfrak{A}}$ returns the same output as $\mathfrak{A}$; by the union bound, this has probability at least $1-2\gamma$. Thus, $\tilde{\mathfrak{A}}$ and $\mathfrak{A}$ have total variation distance at most $2\gamma$.
@@ Appendix C Proofs for statistical sampling
Proof of Lemma 13. When $p\geq\bar{p}$ we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon p)\leq 2e^{-Np\varepsilon^{2}/3}$; for $N$ larger than the stated bound and $p\geq\bar{p}$ this is at most $\gamma/2$. When $p<\bar{p}$, we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon\bar{p})$; by monotonity this is at most $F(N\bar{p},N\varepsilon\bar{p})\leq\gamma/2$.
@@
Proof of Lemma 14. For brevity, let $a=\mu_{\alpha}(x),\hat{a}=\hat{\mu}_{\alpha}(x),\eta=Q(\alpha)e^{-\alpha x},\hat{\eta}=\hat{Q}(\alpha)e^{-\alpha x}$ and $y=\delta/\Delta(x)$. If $a=0$, then $\hat{\pi}(x)=0$ as desired. So suppose $a>0$. For the upper bound, we calculate
@@
Proof of Theorem 15. Consider the following procedure: for each $i=1,\dots,N$ draw $r=\lceil 100\tau\rceil$ copies of $X_{i}$ and compute the sample average $\overline{X_{i}}$. For any pair $i^{\prime}\leq i$, define $Y_{i^{\prime},i}=\prod_{\ell=i^{\prime}}^{i}\overline{X_{\ell}}$.
@@
${\mathbb{E}}[Y^{2}]=\prod_{\ell=i^{\prime}}^{i}(1+\text{Variance}[\overline{X_{\ell}}])=\prod_{\ell=i^{\prime}}^{i}(1+\frac{\text{Variance}[X_{\ell}]}{r})\leq e^{\sum_{\ell=i^{\prime}}^{i}{{\mathbb{S}}[{X_{\ell}}]}/r}\leq e^{(\tau\varepsilon^{2})/r}\leq e^{\varepsilon^{2}/100}$
It now suffices to show $|Y-1|\leq(1-e^{-\varepsilon/2})$. By Chebyshev’s inequality, ${{\mathbb{P}}\!\bm{\left[}|Y-1|>(1-e^{-\varepsilon/2})\bm{}}\leq\frac{\text{Variance}[Y]}{(1-e^{-\varepsilon/2})^{2}}\leq\frac{e^{\varepsilon^{-2}/100}-1}{(1-e^{-\varepsilon/2})^{2}}$. Straightforward calculations show that this is at most $0.07$ for $\varepsilon\in(0,1)$.
@@
We can execute this procedure for $k=O(\log\tfrac{1}{\gamma})$ independent trials and return

### v11

@@ 8.2 Bounds for ${P_{\tt count-weak}}$ in the general integer setting
Let $m=\lfloor n/2\rfloor$. In this construction, we take $d=m$ and $\beta_{\min}=0,\beta_{\max}=m\log 2$. We define $c$ by:
@@
Intuitively, for each $i=1,\dots,m$, there is a small count $c_{2i-1}$ hidden between two larger counts $c_{2i-2},c_{2i}$. If we sample from $\mu_{\beta}$ for $\beta\gg i\log 2$, then $\mu_{\beta}(2i)$ will drown out $\mu_{\beta}(2i-1)$, and likewise if $\beta\ll i\log 2$ then $\mu_{\beta}(2i-2)$ will drown out $\mu_{\beta}(2i-1)$. So the only way to precisely determine each $c_{2i-1}$ is to take roughly $\frac{1}{\delta\varepsilon^{2}}$ samples of $\mu_{\beta}$ with $\beta\approx i\log 2$. Overall, solving ${P^{\delta,\varepsilon}_{\tt count}}$ requires roughly $\Omega(\frac{m}{\delta\varepsilon^{2}})$ samples in total.
@@
Suppose $\delta<1/100$. Then for any $k\in\mathbb{Z}$, we have $Z(k\log 2)\leq 2^{k^{2}+2}$.
@@
For $n>n_{0},q>q_{0},\delta<1/100,\varepsilon<1/100$, where $n_{0},q_{0}$ are absolute constants, the instance can be constructed so that ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{\min\{\sqrt{q},n\}\log(1/\gamma)}{\delta\varepsilon^{2}})$.
@@
Here $\kappa=3\varepsilon$. By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, we may suppose that $n<\sqrt{q}/10$. In this case, we apply Corollary 46(b), where $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}/4+6\varepsilon\leq q$ for $q_{0}$ sufficiently large. For each instance $c^{(\pm i)}$, we use witnesses $x=0,y=2i-1$, where $\Delta(x),\Delta(y)\geq 2\delta\geq\delta e^{2\kappa}$ for $\delta<1/100$, and $|\lambda_{x}^{(i)}-\lambda_{y}^{(i)}|=3\varepsilon$. So ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})\ =\Omega\bigl{(}\frac{n\log(1/\gamma)}{\delta\varepsilon^{2}}\bigr{)}$. ∎
@@ 8.3 Bounds for the continuous setting
In the constructions of Sections 8.1 and 8.2, the complexity bounds have a term $\min\{q,n^{2}\}$. We can modify these constructions to yield continuous-setting instances where the lower bound depends only on $q$ irrespective of $n$. Namely, given a family $c^{(i)}:i=0,\pm 1,\dots,\pm d$ of integer-setting distributions on $[\beta_{\min},\beta_{\max}]$, we construct a new family of continous-setting instances $\tilde{c}$ via
$\tilde{c}^{(i)}_{x}=c^{(i)}_{n(x-1)}\qquad\qquad\text{ for each $i,x$.}$
Theorem 55.
Define $\tilde{\beta}_{\min}=\beta_{\min}n,\tilde{\beta}_{\max}=\beta_{\max}n$. The family $\tilde{c}^{(i)}:i=0,\pm 1,\dots,\pm d$ of continuous-setting distributions on $[\tilde{\beta}_{\min},\tilde{\beta}_{\max}]$ has the following properties:
(a) Each distribution $\tilde{c}^{(i)}$ is supported on $x\in[1,2]$, i.e. it has $\tilde{n}=2$.
(b) Any algorithm to solve ${P^{{\varepsilon}}_{\tt ratio-point}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ for the instances $\tilde{c}$ yields a corresponding algorithm to solve the same problem for $c$, with the same complexity.
(c) The instances $\tilde{c}$ have $\tilde{z}^{(i)}(\tilde{\beta}_{\min},\tilde{\beta}_{\max})\leq z^{(i)}(\beta_{\min},\beta_{\max})+n(\beta_{\max}-\beta_{\min})$ for each $i$.
Proof.
Note that $\tilde{\mu}_{\beta n}^{(i)}(1+x/n)=\mu_{\beta}^{(i)}(x)$ for all $\beta$, and $\tilde{Z}^{(i)}(\beta)=\sum_{x}c^{(i)}_{n(x-1)}e^{\beta x}=\sum_{y}c_{y}^{(i)}e^{\beta(1+y/n)}=e^{\beta}Z^{(i)}(\beta/n).$ So $\tilde{Q}^{(i)}(\tilde{\beta}_{\max})=e^{n(\beta_{\max}-\beta_{\min})}Q(\beta_{\max})$. We can simulate oracle calls to $\tilde{c}$ at input $\tilde{\beta}$ with corresponding oracle calls to $c$ at $\beta=\tilde{\beta}/n$. Thus, solving ${P^{{\varepsilon}}_{\tt ratio-point}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ on $\tilde{c}$ is equivalent to solving them on the original instances $c$. ∎
To get the continuous-setting lower bound $\Omega(\frac{q}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{{\varepsilon}}_{\tt ratio-point}}$ and ${P^{\delta,\varepsilon}_{\tt count-weak}}$, for sufficiently large $q$, we construct the problem instances $c$ from Section 8.1 for value $n^{\prime}=\lfloor\sqrt{q}/2\rfloor,q^{\prime}=q/2$. By Lemma 50 and Theorem 55, we have $\tilde{z}^{(i)}(\tilde{\beta}_{\min},\tilde{\beta}_{\max})\leq n^{\prime}(\beta_{\max}-\beta_{\min})+z^{(i)}(\beta_{\min},\beta_{\max})\leq q$.
Similarly, for the continuous-setting lower bound $\Omega(\frac{\sqrt{q}/\delta}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$, for sufficiently large $q$, we construct the problem instances $c$ of Section 8.2 for value $n^{\prime}=\lfloor\sqrt{q}/10\rfloor,q^{\prime}=q/2$. For $m^{\prime}=\lfloor n^{\prime}/2\rfloor$, we have $n(\beta_{\max}-\beta_{\min})+z^{(i)}(\beta_{\min},\beta_{\max})\leq n^{\prime}(m^{\prime}\log 2-0)+q/2\leq q$.
@@ Appendix A Proofs for error bounds of ${P^{\delta,\varepsilon}_{\tt count}}$
Proof of Theorem 5. Note that $\mu_{\alpha}(x)=\pi(x)e^{\alpha x}/Q(\alpha)$. If $\pi(x)=0$, then the guarantee of ${P^{\delta,\varepsilon}_{\tt count}}$ directly gives $\hat{\mu}_{\alpha}(x)=0=\mu_{\alpha}(x)$. Otherwise, we can use the guarantees of ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$ to show the upper bound on $\hat{\mu}_{\alpha}(x)$:
@@
$\leq e^{\varepsilon}\mu_{\alpha}(x)(1+\varepsilon(1+\delta/\mu_{\alpha}(x)))\qquad\qquad\qquad\qquad\qquad(\text{$\mu_{\alpha}(x)\leq\Delta(x)$ by definition})$
$=e^{\varepsilon}(1+\varepsilon)\mu_{\alpha}(x)+e^{\varepsilon}\delta\leq\mu_{\alpha}(x)+3\varepsilon(\mu_{\alpha}(x)+\delta)\qquad\ (\text{straightforward calculus for $\varepsilon<1/2$})$
@@ Appendix B Correctness with approximate oracles
Here, we define the total variation distance $||\cdot||_{TV}$ by $||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}=\tfrac{1}{2}\sum_{x}|\tilde{\mu}_{\beta}(x)-\mu_{\beta}(x)|$.
@@
Suppose that sampling procedure $\mathfrak{A}$ has cost $T$ and that $||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}\leq\rho$ for a known parameter $\rho$. Let $\tilde{\mathfrak{A}}$ be the algorithm obtained from $\mathfrak{A}$ as follows: (i) replace calls $x\sim\mu_{\beta}$ with calls $x\sim\tilde{\mu}_{\beta}$; (ii) terminate the algorithm after $1/\rho$ steps and return arbitrary answer.
Then $\tilde{\mathfrak{A}}$ has cost $O(T)$, and the total variation distance of the outputs of $\mathfrak{A}$ and $\mathfrak{A}^{\prime}$ is at most $2\rho T$.
@@
The distributions $\mu_{\beta}$ and $\tilde{\mu}_{\beta}$ can be maximally coupled so that samples $x\sim\mu_{\beta}$ and $x\sim\tilde{\mu}_{\beta}$ are identical with probability at least $1-\rho$. Assume the $k^{\text{th}}$ call to $\mu_{\beta}$ in $\mathfrak{A}$ is coupled with the $k^{\text{th}}$ call to $\tilde{\mu}_{\tilde{\beta}}$ in $\tilde{\mathfrak{A}}$ when both calls are defined and $\beta=\tilde{\beta}$. We say the $k^{\text{th}}$ call is good if either (i) both calls are defined and the produced samples are identical, or (ii) $\mathfrak{A}$ has terminated earlier. Note that ${\mathbb{P}}[\mbox{$k^{\text{th}}$ call is good}\>|\>\mbox{all previous calls were good}]\geq 1-\rho$, since the conditioning event implies $\beta=\tilde{\beta}$ (assuming the calls are defined).
@@ 2. $\mathcal{E}_{2}$:
    The number of oracle calls by $\mathfrak{A}$ is at most $1/\rho$. By Markov’s inequality, this has probability at least $1-\frac{{\mathbb{E}}[A]}{1/\rho}\geq 1-\rho T$.
If these two events occur, then $\tilde{\mathfrak{A}}$ returns the same output as $\mathfrak{A}$; by the union bound, this has probability at least $1-2\rho T$. Thus, $\tilde{\mathfrak{A}}$ and $\mathfrak{A}$ have total variation distance at most $2\rho T$.
@@ Appendix C Proofs for statistical sampling
Proof of Lemma 13. When $p\geq\bar{p}$ we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon p)\leq 2e^{-Np\varepsilon^{2}/3}$; for $N$ larger than the stated bound and $p\geq\bar{p}$ this is at most $\gamma/2$. When $p<\bar{p}$, we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon\bar{p})$; by monotonicity this is at most $F(N\bar{p},N\varepsilon\bar{p})\leq\gamma/2$.
@@
Proof of Lemma 14. For brevity, let $a=\mu_{\alpha}(x),\hat{a}=\hat{\mu}_{\alpha}(x),\eta=Q(\alpha)e^{-\alpha x},\hat{\eta}=\hat{Q}(\alpha)e^{-\alpha x}$ and $y=\delta/\Delta(x)$. If $a=0$, then $\hat{a}=0$ since it is an empirical estimate and hence $\hat{\pi}(x)=0$ as desired. So suppose $a>0$. For the upper bound, we calculate
@@
Proof of Theorem 15. As a starting point, consider the following procedure: for each $i=1,\dots,N$ draw $r=\lceil 100\tau\rceil$ copies of $X_{i}$ and compute the sample average $\overline{X_{i}}$; then, for any pair $i^{\prime}\leq i$, set $Y_{i^{\prime},i}=\prod_{\ell=i^{\prime}}^{i}\overline{X_{\ell}}$. We will show that, with high probability, all values $Y_{1,i}$ are close $\mathbb{E}[Y_{1,i}]$.
@@
${\mathbb{E}}[Y^{2}]=\prod_{j=i^{\prime}}^{i}(1+\frac{\text{Variance}[X_{j}]}{r})\leq e^{\sum_{j=i^{\prime}}^{i}{{\mathbb{S}}[{X_{j}}]}/r}\leq e^{(\tau\varepsilon^{2})/r}\leq e^{\varepsilon^{2}/100}$
Now by Chebyshev’s inequality, we have ${{\mathbb{P}}\!\bm{\left[}Y>\notin[e^{-\varepsilon/2},e^{\varepsilon/2}]\bm{}}\leq{{\mathbb{P}}\!\bm{\left[}|Y-1|>(1-e^{-\varepsilon/2})\bm{}}\leq\frac{\text{Variance}[Y]}{(1-e^{-\varepsilon/2})^{2}}\leq\frac{e^{\varepsilon^{-2}/100}-1}{(1-e^{-\varepsilon/2})^{2}}$. Straightforward calculations show that this is at most $0.07$ for $\varepsilon\in(0,1)$.
@@
To obtain the algorithm EstimateProducts, we can run the above procedure for $k=O(\log\tfrac{1}{\gamma})$ independent trials and return